# Beyond the Limits: A Survey of Techniques to Extend the Context Length in Large Language Models

### v2

@@ 1 Introduction
In the rapidly evolving domain of natural language processing (NLP), large language models (LLMs), such as GPT-3, PaLM and LLaMA, emerged as pivotal tools that have proved proficiency in understanding and generating human language including tasks such as language understanding, language generation, complex reasoning and other domains such as computer vision and autonomous driving Brown et al. (2020); Touvron et al. (2023); Chowdhery et al. (2023). In many real-world scenarios, such as multi-turn conversations and document summarization, LLMs are required to comprehend and produce long sequences in order to perform the task accurately during the inference phase. These context sequences are often substantially longer than those the LLMs were trained with, emphasising the fact that LLMs must have the capability to deal with lengthy sequences.
Processing long sequences by LLMs is a non-trivial task, which involves computational, structural, and practical challenges. Notably, increased sequence lengths can exponentially escalate processing requirements, particularly in transformer-based models with self-attention mechanisms. This not only increases the computational cost but also, the memory demands often surpass the capacity of advanced GPUs and thus, impeding efficient training Dao et al. (2022). Hence, the efficiency of attention mechanisms, pivotal in addressing longer sequences, remains a key area of research, aiming to balance computational efficiency with model performance Gu and Dao (2023). Moreover, maintaining contextual understanding and coherence over extended input spans further complicates the scenario, as it requires advanced methods to capture and utilize long-range dependencies. Finally, the evaluation and benchmarking of LLMs on long-sequence tasks also pose a significant challenge, demanding novel metrics and datasets for effective assessment Kwan and Zeng (2023). Altogether, the aforementioned challenges highlight the intricacy and importance of advancing LLMs to proficiently support and utilize long sequences for various tasks.
@@ 2 Length Extrapolation
In this section, we focus on methods whose primary objective is to enable LLMs to effectively support longer sequences. Among these methods, positional extrapolation and interpolation emerge as pivotal methods for extending the model’s capacity to handle sequences longer than those on which the LLM have been originally trained. Furthermore, we explore context window segmentation and sliding, a crucial technique that manipulates input sequences into smaller segments or moves the context window to enable processing of the longer sequence. Lastly, we review the strategy of prompt compression, an innovative approach to condense input prompts efficiently while retaining the essential information.
Positional Extrapolation and Interpolation
@@
Positional extrapolation refers to the model’s ability to handle input sequences that exceed the length of those it was trained on, enabling the preservation of context and coherence over extended sequences. This capability is important for models tasked with understanding and generating lengthy documents or conversations. For example, Attention with Linear Biases (ALiBi) Press et al. (2022) introduces a heuristic of negative causal attention bias, which dispenses with PEs for tokens in the transformer model. ALiBi encodes position information by biasing the query-key attention scores proportionally to the distance between each pair of tokens. As compared to other PE schemes, ALiBi demonstrates superior extrapolation capabilities to unseen sequence lengths. Different from ALiBi, xPOS Sun et al. (2023b) extends causal RoPE, which incorporates a unique exponential decay factor at each dimension of the rotation angle vector, thereby improving length extrapolation. Another approach CLEX Chen et al. (2023a) uses ordinary differential equations to generalize PE scaling. By modeling continuous dynamics with length scaling factors, CLEX effectively overcomes the constraints of traditional positional extrapolation techniques.
@@
On the other hand, positional interpolation deals with the model’s proficiency in inserting or integrating new information within existing sequences. For example, positional interpolation proposed by Chen et al. Chen et al. (2023b) applies linear scaling on the position indices, effectively aligning the maximum position index to correspond with the context window limit previously established during the pre-training phase. Experimental observations indicate that this strategy exhibits greater stability and necessitates fewer fine-tuning steps compared to direct extrapolation methods. Additionally, YaRN Peng et al. (2023b) extends RoPE by adopting an uneven interpolation of frequencies, specifically preserving the high-frequency componentsThis approach avoids losing important positional details that enhances the ability of the model to maintain critical positional information.
Context Window Segmentation and Sliding
@@
LLMs based on transformers are inherently constrained by limited context windows, rendering them incapable of directly integrating or utilizing the entirety of information in long sequences. To mitigate this limitation, various methodologies have been developed to divide the input into segments and apply a sliding window approach to manage the context. One such approach is structured prompting  Hao et al. (2022), which groups demonstration examples and encodes them individually with well-designed position encoding. These encoded examples are then collectively attended to by the test example through a re-scaled attention mechanism, ensuring that each segment receives adequate focus and relevance. Building on the idea of segmenting input, Ratner et al. Ratner et al. (2023) introduces a parallel context window (PCW) that segments the long context into chunks, which segments the long context into chunks and restricts the attention mechanism to operate exclusively within each window. By redeploying positional encoding across these windows, this method ensures efficient processing of long sequences without overwhelming the attention mechanism. Another innovative approach is StreamingLLM Xiao et al. (2023), which addresses the “attention sink” phenomenon. This phenomenon occurs when a significant portion of the attention score is allocated to the initial tokens, regardless of their relevance. StreamingLLM merges window context with the first token, which enables the LLMs trained with a finite-length attention window to be effectively generalized to infinite sequence lengths without requiring additional fine-tuning.
Prompt Compression
@@
Prompt compression refers to methods that shorten original prompts while keeping the important information. This process involves either condensing extensive prompt inputs or learning concise representations of prompts. LLMLingua Jiang et al. (2023) employs streamlined and proficient language models, such as GPT-2 small or LLaMA-7B, to identify and eliminate extraneous tokens within prompts. This method facilitates the efficient execution of inferences with expansive language models, achieving a compression ratio of up to 20 times while maintaining performance with minimal decline. Building on this approach, LongLLMLingua Jiang and Wu (2023) addresses the inherent “lost in the middle” issue observed in LLMs, enhancing the processing of long-context information. This method not only reduces costs but also improves efficiency through prompt compression, resulting in a significant improvement of up to 21.4% in retrieval-augmented generation performance while using only a quarter of the tokens. Further advancing the field, Li et al. Li et al. (2023b) introduces a novel method called “Selective Context”. This approach systematically identifies and prunes redundancy within the input context to streamline the input, making it more compact and optimizing the overall efficiency of language model inferences. MemGPT Packer et al. (2023) is then proposed to overcome the limitations of fixed-length context windows in traditional LLMs. The primary goal is to simulate an infinite context while still effectively utilizing fixed-context models. MemGPT achieves this by autonomously managing its own memory through “function calls” allowing for dynamic context modifications during a task. It establishes a memory hierarchy, akin to traditional operation systems, and treats context windows as constrained memory resources. By enabling the LLM to control its context, MemGPT provides an illusion of longer context length.
@@ 3 Attention Approximation
The foundation of attention approximation lies in the ambition to reduce the computation and memory complexities of vanilla self-attention Vaswani et al. (2017), which increase quadratically with respect to the sequence length $n$, i.e., $O(n^{2})$. This can be achieved by approximating the full-rank attention map with a low-rank counterpart, exploiting the sparse patterns in the attention layers, or simplifying the softmax-related complexity of vanilla attention. These techniques aim to provide efficient approximations that maintain the effectiveness of the attention mechanism while managing long sequences more efficiently.
Low-rank Decomposition
@@
Among the most straightforward yet practical instances of sparse patterns, Block-wise Self Attention Qiu et al. (2020), stands out as an illustrative demonstrations. This method reduces the computation and memory cost by chunking the input sequence into fixed blocks. An alternative strategy involves having an individual token attend to tokens at regular, fixed internals. For instance, Longformer Beltagy et al. (2020) is a sparsifying mechanism that utilizes dilated windows of tokens to construct the attention matrix. LogSparse Li et al. (2019) is another method that sparsifies the attention matrix by restricting consideration to a limited window of tokens, where the window is defined by exponential steps from the token itself. This approach ensures a targeted focus range for each individual token. By employing LogSparse, it is guaranteed that any pair of tokens can exchange attention information with each other, while the memory usage of the transformer can be reduced to $O(n(Log~{}n)^{2})$. LongNet Ding et al. (2023) introduces dilated attention, in which attention allocation decreases exponentially as the distance between tokens increases. This approach exploits mixed dilated rates to accommodate both local and global dependencies between different tokens. By utilizing LongNet, a linear computation complexity, $O(n)$, and a logarithm dependency between tokens can be achieved.
@@
Softmax-free Attention
The efficacy of vanilla attention is often attributed to the softmax operation, which is important for capturing long dependencies. However, this operation causes quadratic complexity in both time and space, impeding the scalability of transformers for long sequences. Replacing the softmax operation can reduce computational complexity, enhancing the efficiency of processing long sequences. This category of approaches is called softmax-free attention.
CosFormer Qin et al. (2022) emulates softmax behaviors through a linear operatorthat re-weights the cosine-based distance. SOFT Lu et al. (2021) employs a Gaussian kernel function to replace the softmax,while SIMA Koohpayegani et al. (2024) opts for normalizing query and key matrices using a simple L1-norm. Another set of approaches replaces softmax with the ReLU function for normalization, demonstrating that this substitution maintains performance, while preserving linear scalability Shen et al. (2023). An alternative class of architectures centers around generalized kernelziable attention, wherein the conventional attention mechanism is formulated as a specific kernel function. For instance, Performer Choromanski et al. (2020) is an approach leveraging positive orthogonal random features to effectively model the attention mechanism into simplifies softmax-free architecture with linear space and time complexity.
Another recently-developed transformer architecture that can be studied under this category (to varying degrees) is RetNet Sun et al. (2023c), which replaces the softmax operation with a D-matrix followed by GroupNorm. The D-matrix introduces exponential decay weighting of previous tokens, diminishing the impact of distant tokens. The incorporation of GroupNorm adds non-linearity, a characteristic once inherent in softmax. A distinguished feature of RetNet is that it can be implemented in both parallel and sequential manners. Accordingly, it can exploit the accelerated token generation during inference, similar to Recurrent Neural Networks (RNN), and exploit the efficiency of parallelization during training.
@@ 4 Attention-free Transformers
State Space Model
@@
SSM is a statistical sequence-to-sequence (seq2seq) model that employs linear projections of hidden states to compute the output sequence based on an input sequence. SSM introduces an RNN-like seq2seq model without non-linearity, which empowers parallel trainingand optimize the inference efficiency. The seq2seq operation based on the states can be analytically unrolled, resembling a convolutional operation with a parametrized kernel. Theoretically, similar to RNN, this convolution operation can extend to infinite length, enabling the computation of outputs without calculating individual states. During the training phase with the entire input sequence, this convolution process can be exceptionally rapid and parallel, setting it apart from the traditional RNN. However, the computational expense of the convolution kernel limits SSM’s application in deep learning until the advent of Structured State Space (S4) Gu et al. (2021). S4 integrates SSM, HIPPO, and structured matrices to solve the complexity of the convolution kernel. HIPPO is a particular representation of the original SSM; takes input states and maps them to higher-dimensional states that can be seen as an online compression of history. However, with a finite size of states, this method cannot remember the entire input. This necessitates the introduction of exponential decay, particularly beneficial for recent past accuracy. Some approaches employ decay techniques, showcasing performance improvements Orvieto et al. (2023); Ma et al. (2022). H3 incorporates two SSMs, enabling local token attention and global token recall through a multiplicative gate mechanism, akin to LSTM gating Fu et al. (2022). Hyena Poli et al. (2023), similar to H3, replaces the S4 layer with an MLP-parameterized global convolution, effectively narrowing the quality gap with the vanilla attention mechanism at scale and achieving comparable perplexity with a reduced computational cost. Mamba Gu and Dao (2023) enhances SSM models by incorporating H3 with MLP, refining reasoning capabilities through a strategic reorganization of the gating mechanism.
Position-dependent Attention
@@
Within this distinct category, a unique form of dependency calculation emerges, where dependencies rely on the position of tokens rather than interactions between them. The Attention-free Transformer (AFT)  Zhai et al. (2021), inspired by attention-based transformers, exclusively employs $\mathbf{K}$ and $\mathbf{V}$ while eliminating $\mathbf{Q}$ and its dot product with $\mathbf{K}$. Instead, AFT introduces a novel learnable matrix $\mathbf{W}$, which acts as a fixed attention map (static routing) consistent across all input sequences. Unlike adaptive weighting in vanilla attention, $\mathbf{W}$ considers only pairwise token positions, disregarding semantic dependencies. To enhance customization based on current input data, key vector ($\mathbf{K}$) accompanies $\mathbf{W}$.
@@ 5 Model Compression
The PTQ approaches can be categorized into weight-only quantization, which only focuses on quantizing the weights and weight-activation quantization, which quantizes both weights and activations. LLM.int8() Dettmers et al. (2022) is the first multi-billion-scale Int8 quantization procedure that reduces memory usage by half during inference, while it maintains the performance the same as that in the full-precision model. GPTQ Frantar et al. (2022) proposes a layer-wise quantization technique, which can further reduce the precision to 3 or 4 bits per weight element, with negligible accuracy degradation. Furthermore, the authors in Lin et al. (2023) find that weights do not carry equal importance for the performance, and accordingly, the quantization error can be significantly reduced by maintaining only 1% of salient weights in full-precision. They propose Activation-aware Quantization (AWQ) method, which retains the weights corresponding to large activations in full-precision. In order to address the significant quantization error resulted from the outliers in activations distribution, Lee et al. (2023) proposes a mixed-precision quantization approach, namely, outlier-aware weight quantization (OWQ), which applies higher precision to the weights associated with outlier activations.
Pruning
@@
Structured pruning aims at removing higher-granularity structures, such as entire neurons, layers, or rows/columns of weight matrices, which can result in a model that retains its original structure but with fewer number of parameters. LLM-Pruner proposed in Ma et al. (2023) is a structural task-agnostic pruning approach that selectively removes non-critical connection structures considering both first-order information and an approximated Hessian information gradient information. As an alternative, Sheared LLaMA Xia et al. (2023) utilizes a two-stage approach for pruning an LLM. In the first stage, it exploits targeted structured pruning to prune a large model to a targeted shape by pruning layers, heads, and intermediate connections. In the second stage, the batches of data are loaded dynamically and the model structure is modified in each training iteration based on losses in various domains. As a result, Sheared LLaMA achieves a compressed model that can outperform the LLMs, with the same size but trained from scratch.
@@
Unstructured pruning involves with pruning individual parameters of a model independently based on their magnitudes or importance, resulting in an irregular sparse structure. Due to the irregularity in the structure and in the memory access patterns, unstructured pruning hinders the efficiency gain that might be achieved through structured pruning, and it requires specialized software and/or hardware for efficient deployment. SparseGPT Frantar and Alistarh (2023) compresses LLMs with billions of parameter by as much as 60%, almost without affecting the performance of the models. However, SparseGPT heavily relies on weight updates. In order to address this issue, the authors in Sun et al. (2023a) propose Wanda that prunes the weights according to novel criterion, which is mainly based on product value of the weights and their input activations.
Multi-query and Group Attention
@@
While multi-head attention has demonstrated its effectiveness in characterizing the correlations among tokens, it suffers from the incremental memory bandwidth cost and longer latency during inference due to repeatedly loading the large $\mathbf{KV}$ tensors as the input sequence length increases. Multi-query attention (MQA) Shazeer (2019) is one of the approaches that address the aforementioned issue. In particular, MQA essentially reuses the same $\mathbf{KV}$ tensors across all attention heads of each query to reduce the memory bandwidth requirements during decoding and thus, allows longer sequences and faster decoding. Given its demonstrated performance with minor quality degradation, MQA has been adopted in several works. Google Chowdhery et al. (2023) trained a LLM named Pathways Language Model (PaLM) with the adoption of MQA to improve decoding speed and later PaLM-2 Anil et al. (2023) is released with improved computation efficiency. Pope et al. (2023) proposed a partition-optimized model that enables up to 32×32$\times$32 × larger context lengths with the help of MQA on LLMs. de Jong et al. (2022) adopts MQA to reduce the cross-attention computation at the decoders in Fusion-in-Decoder (FiD) models with faster inference. More recently, Raymond et al. Li et al. (2023a) introduced StarCoder, a Code LLM, with fast large-batch inference enabled by MQA. The shared $\mathbf{KV}$ tensors idea in MQA also inspired the emergence of other attention computation schemes. In Ainslie et al. (2023), a grouped-query attention (GQA) mechanism is proposed to trade-off performance degradation and speed by sharing a subset of $\mathbf{KV}$ tensors.

### v3

@@ 1 Introduction
In the rapidly evolving domain of natural language processing (NLP), large language models (LLMs), such as GPT-3, PaLM and LLaMA, emerged as pivotal tools that have proved proficiency in understanding and generating human language including tasks such as language understanding, language generation, complex reasoning and other domains such as computer vision and autonomous driving Brown et al. (2020); Touvron et al. (2023); Chowdhery et al. (2024); Wang et al. (2023). In many real-world scenarios, such as multi-turn conversations and document summarization, LLMs are required to comprehend and produce long sequences in order to perform the task accurately during the inference phase. These context sequences are often substantially longer than those the LLMs were trained with, emphasising the fact that LLMs must have the capability to deal with lengthy sequences.
Processing long sequences by LLMs is a non-trivial task, which involves computational, structural, and practical challenges. Notably, increased sequence lengths can exponentially escalate processing requirements, particularly in transformer-based models with self-attention mechanisms. This not only increases the computational cost but also, the memory demands often surpass the capacity of advanced GPUs and thus, impeding efficient training Dao et al. (2022). Hence, the efficiency of attention mechanisms, pivotal in addressing longer sequences, remains a key area of research, aiming to balance computational efficiency with model performance Gu and Dao (2023). Moreover, maintaining contextual understanding and coherence over extended input spans further complicates the scenario, as it requires advanced methods to capture and utilize long-range dependencies. Finally, the evaluation and benchmarking of LLMs on long-sequence tasks also pose a significant challenge, demanding novel metrics and datasets for effective assessment Kwan et al. (2023). Altogether, the aforementioned challenges highlight the intricacy and importance of advancing LLMs to proficiently support and utilize long sequences for various tasks.
@@ 2 Length Extrapolation
In this section, we focus on methods whose primary objective is to enable LLMs to effectively support longer sequences. Among these methods, positional extrapolation and interpolation emerge as pivotal methods for extending the model’s capacity to handle sequences longer than those on which the LLMs have been originally trained. Furthermore, we explore context window segmentation and sliding, a crucial technique that manipulates input sequences into smaller segments or moves the context window to enable processing of the longer sequence. Lastly, we review the strategy of prompt compression, an innovative approach to condense input prompts efficiently while retaining the essential information.
Positional Extrapolation and Interpolation.
@@
Positional extrapolation refers to the model’s ability to handle input sequences that exceed the length of those it was trained on, enabling the preservation of context and coherence over extended sequences. This capability is important for models tasked with understanding and generating lengthy documents or conversations. For example, Attention with Linear Biases (ALiBi) Press et al. (2022) introduces a heuristic of negative causal attention bias, which dispenses with PEs for tokens in the transformer model. ALiBi encodes position information by biasing the query-key attention scores proportionally to the distance between each pair of tokens. As compared to other PE schemes, ALiBi demonstrates superior extrapolation capabilities to unseen sequence lengths. Different from ALiBi, xPOS Sun et al. (2023c) extends causal RoPE, which incorporates a unique exponential decay factor at each dimension of the rotation angle vector, thereby improving length extrapolation. Another approach CLEX Chen et al. (2024) uses ordinary differential equations to generalize PE scaling. By modeling continuous dynamics with length scaling factors, CLEX effectively overcomes the constraints of traditional positional extrapolation techniques.
@@
On the other hand, positional interpolation deals with the model’s proficiency in inserting or integrating new information within existing sequences. For example, positional interpolation proposed by Chen et al. Chen et al. (2023) applies linear scaling on the position indices, effectively aligning the maximum position index to correspond with the context window limit previously established during the pre-training phase. Experimental observations indicate that this strategy exhibits greater stability and necessitates fewer fine-tuning steps compared to direct extrapolation methods. Additionally, YaRN Peng et al. (2023b) extends RoPE by adopting an uneven interpolation of frequencies, specifically preserving the high-frequency components. This approach avoids losing important positional details that enhances the ability of the model to maintain critical positional information.
Context Window Segmentation and Sliding.
@@
LLMs based on transformers are inherently constrained by limited context windows, rendering them incapable of directly integrating or utilizing the entirety of information in long sequences. To mitigate this limitation, various methodologies have been developed to divide the input into segments and apply a sliding window approach to manage the context. One such approach is structured prompting Hao et al. (2022), which groups demonstration examples and encodes them individually with well-designed position encoding. These encoded examples are then collectively attended to by the test example through a re-scaled attention mechanism, ensuring that each segment receives adequate focus and relevance. Building on the idea of segmenting input, Ratner et al. Ratner et al. (2023) introduces a parallel context window (PCW), which segments the long-context into chunks and restricts the attention mechanism to operate exclusively within each window. By redeploying positional encoding across these windows, this method ensures efficient processing of long sequences without overwhelming the attention mechanism. Another innovative approach is StreamingLLM Xiao et al. (2024), which addresses the “attention sink” phenomenon. This phenomenon occurs when a significant portion of the attention score is allocated to the initial tokens, regardless of their relevance. StreamingLLM merges window context with the first token, which enables the LLMs trained with a finite-length attention window to be effectively generalized to infinite sequence lengths without requiring additional fine-tuning.
Prompt Compression.
@@
Prompt compression refers to methods that shorten original prompts while keeping the important information. This process involves either condensing extensive prompt inputs or learning concise representations of prompts. LLMLingua Jiang et al. (2023a) employs streamlined and proficient language models, such as GPT-2 small or LLaMA-7B, to identify and eliminate extraneous tokens within prompts. This method facilitates the efficient execution of inferences with expansive language models, achieving a compression ratio of up to 20 times while maintaining performance with minimal decline. Building on this approach, LongLLMLingua Jiang et al. (2023b) addresses the inherent “lost in the middle” issue observed in LLMs, enhancing the processing of long-context information. This method not only reduces costs but also improves efficiency through prompt compression, resulting in a significant improvement of up to 21.4% in retrieval-augmented generation performance while using only a quarter of the tokens. Further advancing the field, Li et al. Li et al. (2023b) introduce a novel method called “Selective Context”. This approach systematically identifies and prunes redundancy within the input context to streamline the input, making it more compact and optimizing the overall efficiency of language model inferences. MemGPT Packer et al. (2023) is then proposed to overcome the limitations of fixed-length context windows in traditional LLMs. The primary goal is to simulate an infinite context while still effectively utilizing fixed-context models. MemGPT achieves this by autonomously managing its own memory through “function calls” allowing for dynamic context modifications during a task. It establishes a memory hierarchy, akin to traditional operation systems, and treats context windows as constrained memory resources. By enabling the LLM to control its context, MemGPT provides an illusion of longer context length.
@@ 3 Attention Approximation
The foundation of attention approximation lies in the ambition to reduce the computation and memory complexities of vanilla self-attention Vaswani et al. (2017), which increases quadratically with respect to the sequence length $n$, i.e., $O(n^{2})$. This can be achieved by approximating the full-rank attention map with a low-rank counterpart, exploiting the sparse patterns in the attention layers, or simplifying the softmax-related complexity of vanilla attention. These techniques aim to provide efficient approximations that maintain the effectiveness of the attention mechanism while managing long sequences more efficiently.
Low-rank Decomposition.
@@
Among the most straightforward yet practical instances of sparse patterns, Block-wise Self Attention Qiu et al. (2020), stands out as an illustrative demonstrations. This method reduces the computation and memory cost by chunking the input sequence into fixed blocks. An alternative strategy involves having an individual token attend to tokens at regular, fixed internals. For instance, Longformer Beltagy et al. (2020) is a sparsifying mechanism that utilizes dilated windows of tokens to construct the attention matrix. LogSparse Li et al. (2019) is another method that sparsifies the attention matrix by restricting consideration to a limited window of tokens, where the window is defined by exponential steps from the token itself. This approach ensures a targeted focus range for each individual token. By employing LogSparse, it is guaranteed that any pair of tokens can exchange attention information with each other, while the memory usage of the transformer can be reduced to $O(n(Log~{}n)^{2})$. LongNet Ding et al. (2023) introduces dilated attention, in which attention allocation decreases exponentially as the distance between tokens increases. This approach exploits mixed dilated rates to accommodate both local and global dependencies between different tokens. It has been shown that by utilizing LongNet, a linear computation complexity, $O(n)$, and a logarithm dependency between tokens can be achieved.
@@
Softmax-free Attention.
The efficacy of vanilla attention Vaswani et al. (2017) is often attributed to the softmax operation, which is important for capturing long dependencies. However, this operation causes quadratic complexity in both time and space, impeding the scalability of transformers for long sequences. Replacing the softmax operation can reduce computational complexity, enhancing the efficiency of processing long sequences. This category of approaches is called softmax-free attention.
CosFormer Qin et al. (2022) emulates softmax behaviors through a linear operatorthat re-weights the cosine-based distance. SOFT Lu et al. (2021) employs a Gaussian kernel function to replace the softmax, while SIMA Koohpayegani and Pirsiavash (2024) opts for normalizing query and key matrices using a simple L1-norm. Another set of approaches replaces softmax with the ReLU function for normalization, demonstrating that this substitution maintains performance, while preserving linear scalability Shen et al. (2023). An alternative class of architectures centers around generalized kernelziable attention, wherein the conventional attention mechanism is formulated as a specific kernel function. For instance, Performer Choromanski et al. (2021) is an approach leveraging positive orthogonal random features to effectively model the attention mechanism into simplified softmax-free architecture with linear space and time complexity.
Another recently-developed transformer architecture that can be studied under this category (to varying degrees) is RetNet Sun et al. (2023b), which replaces the softmax operation with a D-matrix followed by group normalization (GroupNorm). The D-matrix introduces exponential decay weighting of previous tokens, diminishing the impact of distant tokens. The incorporation of GroupNorm adds non-linearity, a characteristic once inherent in softmax. A distinguished feature of RetNet is that it can be implemented in both parallel and sequential manners. Accordingly, it can exploit the accelerated token generation during inference, similar to Recurrent Neural Networks (RNN), and exploit the efficiency of parallelization during training.
@@ 4 Attention-free Transformers
State Space Model.
@@
SSM is a statistical sequence-to-sequence (seq2seq) model that employs linear projections of hidden states to compute the output sequence based on an input sequence. SSM introduces an RNN-like seq2seq model without non-linearity, which empowers parallel trainingand optimize the inference efficiency. The seq2seq operation based on the states can be analytically unrolled, resembling a convolutional operation with a parametrized kernel. Theoretically, similar to RNN, this convolution operation can extend to infinite length, enabling the computation of outputs without calculating individual states. During the training phase with the entire input sequence, this convolution process can be exceptionally rapid and parallel, setting it apart from the traditional RNN. However, the computational expense of the convolution kernel limits SSM’s application in deep learning until the advent of Structured State Space (S4) Gu et al. (2022). S4 integrates SSM, HIPPO Gu et al. (2020), and structured matrices to solve the complexity of the convolution kernel. HIPPOO Gu et al. (2020) is a particular representation of the original SSM, which takes input states and maps them to higher-dimensional states that can be seen as an online compression of history. However, with a finite size of states, this method cannot remember the entire input. This necessitates the introduction of exponential decay, particularly beneficial for recent past accuracy. Some approaches employ decay techniques, showcasing performance improvements Orvieto et al. (2023); Ma et al. (2023b). Hungry Hungry Hippos (H3) Fu et al. (2023) incorporates two SSMs, enabling local token attention and global token recall through a multiplicative gate mechanism, akin to LSTM gating. Hyena Poli et al. (2023), similar to H3, replaces the attention layer by interleaving implicitly parametrized long convolutions and data-controlled gating, effectively narrowing the quality gap with the vanilla attention mechanism at scale and achieving comparable perplexity with a reduced computational cost. Mamba Gu and Dao (2023) enhances SSMs by incorporating H3 with multi-layer perceptrons (MLP), refining reasoning capabilities through a strategic reorganization of the gating mechanism.
Position-dependent Attention.
@@
Within this distinct category, a unique form of dependency calculation emerges, where dependencies rely on the position of tokens rather than interactions between them. The Attention-free Transformer (AFT)  Zhai et al. (2021), inspired by attention-based transformers, exclusively employs $\mathbf{K}$ and $\mathbf{V}$ while eliminating $\mathbf{Q}$ and its dot product with $\mathbf{K}$. Instead, AFT introduces a novel learnable matrix $\mathbf{W}$, which acts as a fixed attention map (static routing) consistent across all input sequences. Unlike adaptive weighting in vanilla attention, $\mathbf{W}$ considers only pairwise token positions, disregarding semantic dependencies. To enhance customization based on current input data, $\mathbf{K}$ accompanies $\mathbf{W}$.
@@ 5 Model Compression
The PTQ approaches can be categorized into weight-only quantization, which only focuses on quantizing the weights and weight-activation quantization, which quantizes both weights and activations. LLM.int8() Dettmers et al. (2022) is the first multi-billion-scale INT8 quantization procedure that reduces memory usage by half during inference, while it maintains the performance the same as that in the full-precision model. OPTQ Frantar et al. (2023) proposes a layer-wise quantization technique, which can further reduce the precision to 3 or 4 bits per weight element, with negligible accuracy degradation. Furthermore, Lin et al. Lin et al. (2024) find that weights do not carry equal importance for the performance, and accordingly, the quantization error can be significantly reduced by maintaining only 1% of salient weights in full-precision. They propose Activation-aware Quantization (AWQ) method, which retains the weights corresponding to large activations in full-precision. In order to address the significant quantization error resulted from the outliers in activations distribution, Lee et al. Lee et al. (2024) propose a mixed-precision quantization approach, namely, outlier-aware weight quantization (OWQ), which applies higher precision to the weights associated with outlier activations.
Pruning.
@@
Structured pruning aims at removing higher-granularity structures, such as entire neurons, layers, or rows/columns of weight matrices, which can result in a model that retains its original structure but with fewer number of parameters. LLM-Pruner Ma et al. (2023a) is a structural task-agnostic pruning approach that selectively removes non-critical connection structures considering both first-order information and an approximated Hessian information gradient information. As an alternative, Sheared LLaMA Xia et al. (2024) utilizes a two-stage approach for pruning an LLM. In the first stage, it exploits targeted structured pruning to prune a large model to a targeted shape by pruning layers, heads, and intermediate connections. In the second stage, the batches of data are loaded dynamically and the model structure is modified in each training iteration based on losses in various domains. As a result, Sheared LLaMA achieves a compressed model that can outperform the LLMs, with the same size but trained from scratch.
@@
Unstructured pruning involves with pruning individual parameters of a model independently based on their magnitudes or importance, resulting in an irregular sparse structure. Due to the irregularity in the structure and in the memory access patterns, unstructured pruning hinders the efficiency gain that might be achieved through structured pruning, and it requires specialized software and/or hardware for efficient deployment. SparseGPT Frantar and Alistarh (2023) compresses LLMs with billions of parameter by as much as 60%, almost without affecting the performance of the models. However, SparseGPT heavily relies on weight updates. In order to address this issue, Sun et al. Sun et al. (2023a) propose Wanda that prunes the weights according to novel criterion, which is mainly based on product value of the weights and their input activations.
Multi-query and Group Attention.
@@
While multi-head attention has demonstrated its effectiveness in characterizing the correlations among tokens, it suffers from the incremental memory bandwidth cost and longer latency during inference due to repeatedly loading the large $\mathbf{KV}$ tensors as the input sequence length increases. Multi-query attention (MQA) Shazeer (2019) is one of the approaches that address the aforementioned issue. In particular, MQA essentially reuses the same $\mathbf{KV}$ tensors across all attention heads of each query to reduce the memory bandwidth requirements during decoding and thus, allows longer sequences and faster decoding. Given its demonstrated performance with minor quality degradation, MQA has been adopted in several works. Google Chowdhery et al. (2024) trains a LLM named Pathways Language Model (PaLM) with the adoption of MQA to improve decoding speed and later PaLM-2 Anil et al. (2023) is released with improved computation efficiency. Pope et al. Pope et al. (2023) propose a partition-optimized model that enables up to 32×32$\times$32 × larger context lengths with the help of MQA on LLMs. de Jong et al. de Jong et al. (2023) adopts MQA to reduce the cross-attention computation at the decoders in Fusion-in-Decoder models with faster inference. More recently, Li et al. Li et al. (2023a) introduce StarCoder, a Code LLM, with fast large-batch inference enabled by MQA. The shared $\mathbf{KV}$ tensors idea in MQA also inspired the emergence of other attention computation schemes. A grouped-query attention (GQA) Ainslie et al. (2023) mechanism is proposed to trade-off performance degradation and speed by sharing a subset of $\mathbf{KV}$ tensors.